How many grams are in 238 moles of arsenic? The atomic mass of arsenic is 33 g.

1
Expert's answer
2022-03-12T01:58:32-0500

n(As) = 238 moles;

M(As) = 33 g/mole;

n = m/M

m(As) = n(As) * M(As) = 238 * 33 = 7854 g.

Answer: 7854 g.